DVI2TTY(Local) DVI2TTY(Local) NAME dvi2tty−previewadvi−fileonanordinaryasciiterminal SYNOPSIS dvi2tty[options]dvi−file DESCRIPTION dvi2ttyconvertsaTeXDVI−filetoaformatthatisapprporiateforterminalsandlineprinters. TheprogramisintendedtobeusedforpreliminaryproofreadingofTeX-eddocuments.By defaulttheoutputisdirectedtotheterminal,possiblythroughapager(dependingonhowthe programwasinstalled),butitcanbedirectedtoafileorapipe. Theoutputleavesmuchtobedesired,butisstillusefullifyouwanttoavoidwalkingtothelaserprinter(orwhatever)foreachiterationofyourdocument. Sincedvi2ttyproducesoutputforterminalsandlineprinterstherepresentationofdocumentsis naturallyquiteprimitive.Fontchangesaretotallyignored,whichimpliesthatspecialsymbols, suchasmathematicalsymbols,getmappedintothecharactersatthecorrespondingpositionsin the"standard"fonts. Ifthewidthoftheoutputtextrequiresmorecolumnsthanfitsinoneline(c.f.the−woption)it isbrokenintoseverallinesbydvi2ttyalthoughtheywillbeprintedasonelineonregularTeX outputdevices(e.g.laserprinters).Toshowthatabrokenlineisreallyjustonelogicallinean asterisk(‘‘*’’)inthelastpositionmeansthatthelogicallineiscontinuedonthenextphysicalline outputbydvi2tty.Suchacontinuationlineisstartedwithaaspaceandanasteriskinthefirst twocolumns. OptionsmaybespecifiedintheenvironmentvariableDVI2TTY.Anyoptiononthecommandline,conflictingwithoneintheenvironment,willoverridetheonefromtheenvironment. Options: −ofile Writeoutputtofile‘‘file’’. −plist Printthepageschosenbylist.NumbersrefertoTeX−pagenumbers(knownas\count0). Anexampleofformatforlistis‘‘1,3:6,8’’tochoosepages1,3through6and8.Negative numberscanbeusedexactlyasinTeX,eg−1comesbefore−4asin‘‘−p-1:-4,17’’. −Plist Like−pexceptthatpagenumbersrefertothesequentialorderingofthepagesinthe dvi−file.Negativenumbersdon’tmakealotofsensehere... −wnSpecifyterminalwidthn.Legalrange16−132.Defaultis80.Ifyourterminalhasthe abilitytodisplayin132columnsitmightbeagoodideatouse−w132andtoggletheterminalintothismodeasoutputwillprobablylooksomewhatbetter. −qDon’tpipetheoutputthroughapager.Thismaybethedefaultonsomesystems (dependingonthewhimsoftheSAinstallingtheprogram). −fPipethroughapager,$PAGERifdefined,orwhateveryourSAcompiledin(often ‘‘more’’).Thismaybethedefault,butitisstillokaytoredirectoutputwith‘‘>’’,the pagerwillnotbeusedifoutputisnotgoingtoaterminal. −FSpecifythepagerprogramtobeused.Thisoveridesthe$PAGERandthedefaultpager. −Fprog Use‘‘prog’’asprogramtopipeoutputinto.Canbeusedtochooseanalternatepager(eg ‘‘-Fless’’). −lMarkpagebreakswiththetwo-charactersequence‘‘ˆ L’’.Thedefaultistomarkthemwith aformfeedcharacter. 7June1986 1 DVI2TTY(Local) DVI2TTY(Local) −uDon’tmakeanyattemptstofindspecialScandinaviancharacters.Ifsuchcharactersare inthetexttheywillmapto‘‘a’’and‘‘o’’.ThisisprobablythedefaultoutsideofScandinavia.(TheSAmadethedecisionwhentheprogramwasinstalled.) −sTrytofindthespecialScandinaviancharactersthatonmost(?)terminalsinScandinavia aremappedto‘‘{|}[\]’’.Thiscanbethedefault,andoutputfromfilesnotcontaining thesespecialcharacterswillbeidenticalregardlessofthisoption. FILES /usr/ucb/moreprobablythedefaultpager. ENVIRONMENT PAGERthepagertouse. DVI2TTYcanbesettoholdcommandlineoptions. SEEALSO TeX,dvi2ps AUTHOR SvanteLindahl,RoyalInstituteofTechnology,Stockholm ImprovedCversion:MarcelMol {seismo,mcvax}!enea!ttds!zap marcel@duteca.UUCP BUGS Blanksbetweenwordsgetlostquiteeasy.Thisislesslikelyifyouareusingawideroutputthan thedefault80. Onlyonefilemaybespecifiedonthecommandline. 7June1986 2